The following items appeared on the Balance Sheet of Bellaire Inc. on 31 December 2020. All...

  1. The following items appeared on the Balance Sheet of Bellaire Inc. on 31 December 2020.

All amounts in $

2020

2019

Trade Receivables (Gross)

880,100

765,000

Allowance for Doubtful Debts

(?)

(52,800)

Net Trade Receivables

?

827,300

In addition, the Ageing analysis for the company was as follows:

Age

Trade Rec.

Default rate

0-30 days

510,000

2%

31-60 days

205,000

6%

61-90 days

120,000

15%

90+ days

45,100

50%

Using the above information, compute the Allowance for Doubtful Debts and the Net Trade Receivables for the company for 31.12.2020. Assuming the company recorded $35,000 of write-offs during 2020, what was the amount of bad debt expense that must have been recorded during 2020? Please show your calculations to receive partial credit. [easy 10 points]
Answer #1

Estimated ending allowance for doubtful accounts

Age Trade rec Default rate $ uncollectible
0-30 days 510000 2% 10200
31-60 days 205000 6% 12300
61-90 days 120000 15% 18000
90+ days 45100 50% 22550
Total 63050

Allowance for doubtful account after write off = 52800-35000 = 17800

Bad debt expense = 63050-17800 = 45250